OpenStudy (anonymous):

can ANYONE help me:/ 5x+4y=6 -2x-3y=-1

OpenStudy (anonymous):

simplifying linear equations!

OpenStudy (radar):

You have a "system" of linear equations. Are you wanting a solution for x and y?

OpenStudy (anonymous):

yes.

OpenStudy (radar):

The two most common methods of solving something like this is the "substitution method" or the "elimination" method, which method you prefer?

OpenStudy (anonymous):

whichever would be easier, but usually substitution.

OpenStudy (radar):

I think the elimination method is easier, but substitution works ok too. What do you want to substitute the x or y?

OpenStudy (anonymous):

ok then I'd rather do elimination. (:

OpenStudy (anonymous):

in that case, distribute a 2 through 5x+4y=6. distribute a 5 to -2x-3y=-1

OpenStudy (radar):

OK, multiply the first equation (all terms by 2 and follow that by multiplying the second equation by 5 (all terms) doing so, will get you the following: 10x+8y=12 -10x-15y=-5 Did you get that also?

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

ok. so 2(5x+4y=6) would be 10x+8y=12. and 5(-2x-3y=-1) would be -10x-15y=-5

OpenStudy (radar):

Ok add the two equations and you will only have y left, you have eliminated the X -7y=7 y=-1 real simple wasn't it?

OpenStudy (anonymous):

then line em up. 10x-10x cancels out. 8y-15y is -7y and 12-5 is 7. -7y=7, y=-1

OpenStudy (anonymous):

yes! thank you both

OpenStudy (radar):

Now that you have found y, just plug it in either of the two first equations and solve for x.

OpenStudy (anonymous):

so that is 5x=10 and x=2

OpenStudy (anonymous):

hope it helps